# Optimizing Distributed Reinforcement Learning with Reactor Model and Lingua Franca

### 07 Dec 2023 v1

@@ Abstract
Distributed Reinforcement Learning (RL) frameworks are essential for mapping RL workloads to multiple computational resources, allowing for faster generation of samples, estimation of values, and policy improvement. These computational paradigms require a seamless integration of training, serving, and simulation workloads. Existing frameworks, such as Ray, are not managing this orchestration efficiently. In this study, we’ve proposed a solution implementing Reactor Model, which enforces a set of actors to have a fixed communication pattern. This allows the scheduler to eliminate works needed for synchronization, such as acquiring and releasing locks for each actor or sending and processing coordination-related messages. Our framework, Lingua Franca (LF), a coordination language based on the Reactor Model, also provides a unified interface that allows users to automatically generate dataflow graphs for distributed RL. On average, LF outperformed Ray in generating samples from OpenAI Gym and Atari environments by 1.21x and 11.62x, reduced the average training time of synchronized parallel Q-learning by 31.2%, and accelerated Multi-Agent RL inference by 5.12x.
1 Introduction
@@
The field of machine learning (ML) has witnessed an exponential increase in computational requirements for training complex models, particularly supervised learning powered by deep neural networks. This complexity has necessitated the creation of new frameworks focused on training these networks and leveraging specialized hardware to reduce training times. Examples include TensorFlow abadi2016tensorflow , MXNet chen2015mxnet , and PyTorch paszke2019pytorch . Beyond classical supervised learning, emerging AI applications are increasingly required to operate in dynamic environments and pursue long-term goals, problems that reinforcement learning (RL) are well suited for. RL involves learning to behave in uncertain environments based on feedback from interaction with the environment. Optimal behavior can be specified via appropriately chosen reward functions of the data. This approach has already led to significant achievements such AlphaGo silver2016mastering , and more recently, the success of ChatGPT OpenAI_GPT4_2023 . RL applications span various domains, including traffic systems (chen2018auto, ), UAVs (kaufmann2023champion, ), large language models (LLMs) OpenAI_GPT4_2023 , and dexterous manipulation (akkaya2019solving, ).
@@
Deep RL presents new additional challenges beyond traditional RL for distributed frameworks. Deep RL, unlike its traditional counterparts, demands a sequential interaction between the agent and the environment, making it predominantly utilize the CPU (central processing unit) ali2019view . However, the scalability of deep RL models, particularly in handling complex state-action associations, hinges on efficiently leveraging both CPUs and graphical processing units (GPUs) inci2020architectural ; espeholt2019seed . The combination of RL with deep neural networks in deep RL models necessitates a balanced approach in computational resource allocation. The processing speed, especially when updating policies involving millions of parameters, becomes a critical factor. The need to balance CPU and GPU resources, coupled with the limitations of existing frameworks, highlights the need for innovative approaches in developing efficient, scalable, and versatile systems capable of supporting the dynamic and complex nature of modern AI and ML applications. This paper aims to delve into these challenges and explore potential solutions, paving the way for more efficient and effective reinforcement learning systems.
To this end, we introduce an optimized version of Lingua Franca (LF) lohstroh2021toward , a polyglot coordination language for concurrent and time-sensitive applications initially designed for embedded systems, which was tailored to address the unique challenges of RL applications. LF stands out in its ability to effectively handle a diverse range of workloads, from lightweight, stateless computations needed for simulation to the more intensive, long-running computations required for training.
A key feature of LF is its unified interface, which is adept at expressing reactor-based computations (e.g. rollout worker reactors, replay buffer reactors, and learner reactors). This adaptability is crucial for efficiently managing the diverse computational demands of RL applications. The computational paradigm in LF allows for effective load balancing of simulations, handling inputs and state spaces in shared memory, and provides robustness against failures. On the other hand, the reactor-based model is optimized for stateful computations, such as those found in model training, and it facilitates shared mutable state to clients, like in a parameter server architecture.
LF is a distributed framework that successfully unifies training, simulation, and serving. Our framework also provides a unified interface that enables users to easily define abstractions in RL. On average, LF outperformed Ray moritz2018ray in generating samples from OpenAI Gym brockman2016openai and Atari environments by 1.21x and 11.62x, reduced the average training time of synchronized parallel Q-learning by 31.2%, and accelerated Multi-Agent RL inference by 5.12x.
@@
Contributions: In this paper, we demonstrate how the reactor-oriented paradigm implemented in Lingua Franca optimizes utilization of computational resources and efficiently parallelizes RL workloads. We then describe the mechanisms of LF runtime, illustrating how it exploits parallelism and differs from the actor-based framework, Ray. Additionally, our investigation reveals that multithreading offers greater advantages than multiprocessing in distributed RL. We then introduce an optimized Reactor C runtime with customized Python, enabling true parallelism and efficient coordination of threads. We further present an extensive evaluation on the generations of samples from OpenAI Gym Environments, inference of multi-Agent RL, and synchronized parallel actor critic demonstrating that the LF runtime outperforms Ray in the training, inference, and simulation for RL.
2 Distributed Reinforcement Learning and Requirements
@@
Reinforcement Learning (RL) algorithms aim to enhance an agent’s policy performance within a given environment, often represented through a simulator. These algorithms alternate cycles of data collection with the latest policy, value estimation using the latest data (and possibly data collected from previous policies), and policy improvement. Depending on the signal provided by the reward function on the collected data and the efficiency of the algorithm itself, the reinforcement learning process can vary in the amount of data/computational resources needed to run until satisfactory performance is achieved. For most tasks of interest, these data and compute requirements are large, motivating distributed reinforcement learning frameworks that can more efficiently leverage parallel computational paradigms.
@@
Many aspects of the reinforcement learning problem can be parallelized (samsami2020distributed, ). Data collection can be split across multiple worker threads, learning can be split across multiple worker threads with each thread maintaining its own value function and policy parameters, updates to the neural networks can be performed in parallelized fashions using popular deep learning frameworks (abadi2016tensorflow, ; bradbury2018jax, ; paszke2017automatic, ), and simulators can themselves be parallelized, for example on accelerated computing infrastructure (puig2023habitat3, ; szot2021habitat, ; habitat19iccv, ; shacklett2023extensible, ).
2.1 Variants in RL Algorithms
@@
Single-Agent Training: The most fundamental scenario in RL is training a single agent, which involves repeatedly applying the steps of rollout, replay, and optimization. Synchronous algorithms like A2C (mnih2016asynchronous, ) and PPO (schulman2017proximal, ) follow these steps in sequence, whereas asynchronous variants (e.g., A3C (mnih2016asynchronous, ), Ape-X (horgan2018distributed, ), APPO (zeng2020visual, ), IMPALA (espeholt2018impala, )) overlap rollout and optimization steps to enhance data throughput.
@@
Multi-Agent Training: Multi-agent training involves multiple agents interacting within the environment, either cooperatively or competitively. While the dataflow structure resembles that of single-agent training, complexities arise when customizing training for individual agents. For instance, if agents require optimization at different frequencies or are trained with distinct algorithms, the training dataflow must accommodate multiple iterative loops with varying parameters.
@@
Model-Based Algorithms: Model-based RL algorithms aim to learn the transition dynamics of the environment to increase training efficiency (moerland2023model, ). This adds a supervised training component to standard distributed RL, involving training one or more dynamics models with environment-generated data.
@@
Challenges and Opportunities in Framework Optimization: Frameworks like Ray have made strides in CPU/GPU orchestration for RL, but they face efficiency challenges due to the actor model’s inherent communication overhead and synchronization demands. Ideal frameworks for RL should support heterogeneous computations, flexibility in computational models, dynamic execution, and large data handling, while integrating seamlessly with deep learning libraries and simulation frameworks.
@@
We argue that distributed RL frameworks can better handle millions of tasks per second by providing deterministic concurrency. Current frameworks, on the other hand, fall short in efficiently meeting the evolving demands of AI applications, indicating a significant gap in the field.
This research aims to propose a novel system architecture that addresses these comprehensive requirements, bridging the gap in the current landscape and pushing the boundaries of what’s possible in distributed RL.
3 Introduction to Actor and Reactor Model
3.1 Actor Model
@@
The Actor model hewitt2010actor , a foundational concept in concurrent computing, emerged in the 1970s as a response to the increasing complexity and interactivity in computer systems. It introduced a novel way to handle concurrent operations by conceptualizing "actors" as the primary units of computation. These actors are analogous to objects in object-oriented programming (OOP) rentsch1982object , but they are designed specifically to address the challenges of concurrency. Each actor represents a self-contained unit with its own local state, and actors interact with each other exclusively through asynchronous message passing. This model contrasts with traditional approaches that often rely on shared state and synchronization mechanisms like locks, which can lead to issues like deadlocks and race conditions.
@@
Furthermore, the Actor model introduces a flexible approach to message processing, without enforcing any strict order in which messages must be processed. This characteristic is particularly beneficial in distributed systems where message delivery times can vary unpredictably. Actors can process incoming messages in different sequences, and the model does not guarantee that messages will arrive in the order they were sent. This flexibility allows for more efficient utilization of resources and can lead to more robust system designs that are tolerant of delays and variable message delivery times. Overall, the Actor model’s emphasis on independent, concurrent actors and asynchronous communication makes it a powerful paradigm for building scalable and resilient concurrent and distributed systems.
3.2 Reactor Model
@@
The Reactor model lohstroh2020reactors shows great potential as an alternative to the actor model, enabling efficient and deterministic concurrency. The newly proposed Reactor model represents an advancement in deterministic reactive systems, providing a structured framework for creating complex, reactive RL applications. Central to this model are the concepts of "reactors" and "reactions." Reactors can be interpreted as deterministic actors, but instead of responding to messages, they react to discrete events, each linked to a specific logical time, denoted by a "tag." These events can trigger reactions within a reactor, similar to message handlers in actor systems, but with a key difference: reactions in reactors are governed by a defined order, ensuring determinism. Reactions share some similarities with message handlers in traditional actor models but differ in their triggering mechanism. Instead of messages, they are activated by discrete events, which can also be generated by reactions. Each event associates a value with a tag, representing its logical release time within the system. Reactions can access and modify state shared with other reactions in the same reactor, but interaction between different reactors is solely through events. This design choice illustrates the model’s deterministic nature, as the order of reaction executions is predictable and subject to strict constraints, such as tag order and execution order for reactions within the same reactor.
@@
The Reactor model can be thought of as a "sparse synchronous model." This means that synchronous-reactive interactions at a particular logical time may be confined to isolated parts of the system. When a reaction executes, it exclusively accesses the reactor’s state. Moreover, for reactions triggered by the same event (or events with the same tag), their execution order is pre-defined, reinforcing deterministic behavior. Reactors are composed of ports (inputs and outputs), hierarchy, and actions. The term "reactors" not only relate to actors in actor systems but also aligns with the synchronous reactive programming paradigm, prominent in languages like Esterel, Signal, and Lustre. Unlike traditional actors, reactors don’t directly reference their peers. Instead, they use named and typed ports for interconnection, enabling a hierarchical design. This hierarchy, besides facilitating deterministic behavior, also serves as a scoping mechanism for ports and imposes constraints on connection types.
@@
Additionally, reactors feature actions, a variant of ports used for scheduling future events within the same reactor or as a synchronization mechanism between internal logic and asynchronous external events. This design choice provides a bridge between deterministic internal logic and the nondeterministic external world, like sensor data, environment states, or network messages. The Reactor model also incorporates state variables within reactors. The shared resources, like replay buffers, model parameters, environment states in RL, are key motivators for grouping reactions in a single reactor. However, reactors themselves do not share state, ensuring isolation between reactors and allowing parallel execution of reactions in different reactors, unless a connection necessitates sequential execution.
@@
Connections between reactors establish explicit communication channels. These connections reveal dependencies that are crucial for scheduling decisions honoring data dependencies. The Reactor model simplifies the declaration of these dependencies by breaking down functionality into reactions with well-defined lexical scopes, thus eliminating dependencies out of scope. The Reactor model’s execution is governed by a run time environment, which adheres to a discrete-event model to ensure determinacy. This environment is responsible for maintaining a global event queue and a reaction queue, managing logical time, and executing reactions in a predetermined order. Details regarding the scheduler are described in the next section.
3.3 Ray

### 02 Feb 2024 v2

# Efficient Parallel Reinforcement Learning Framework using the Reactor Model

@@ Abstract.
Parallel Reinforcement Learning (RL) frameworks are essential for mapping RL workloads to multiple computational resources, allowing for faster generation of samples, estimation of values, and policy improvement. These computational paradigms require a seamless integration of training, serving, and simulation workloads. Existing frameworks, such as Ray, are not managing this orchestration efficiently, especially in RL tasks that demand intensive input/output and synchronization between actors on a single node. In this study, we have proposed a solution implementing the reactor model, which enforces a set of actors to have a fixed communication pattern. This allows the scheduler to eliminate work needed for synchronization, such as acquiring and releasing locks for each actor or sending and processing coordination-related messages. Our framework, Lingua Franca (LF), a coordination language based on the reactor model, also supports true parallelism in Python and provides a unified interface that allows users to automatically generate dataflow graphs for RL tasks. In comparison to Ray on a single-node multi-core compute platform, LF achieves 1.21x and 11.62x higher simulation throughput in OpenAI Gym and Atari environments, reduces the average training time of synchronized parallel Q-learning by 31.2%, and accelerates multi-agent RL inference by 5.12x.
Parallel Computing, Reinforcement Learning, Programming Languages, Machine Learning, Model of Computation
conference: June 17-21, 2024; Nantes, France; $c$s: Computing methodologies Parallel programming languages
1. Introduction
@@
The field of machine learning (ML) has witnessed an exponential increase in computational requirements for training models, which tend to be increasingly larger deep neural networks. This complexity has necessitated the creation of new frameworks focused on training these networks and leveraging specialized hardware to reduce training times. Examples include TensorFlow (Abadi et al., 2016), MXNet (Chen et al., 2015), PyTorch (Paszke et al., 2019), and Deepspeed (Rasley et al., 2020). Beyond classical supervised learning, emerging AI applications are increasingly required to operate in dynamic environments and pursue long-term goals, problems that reinforcement learning (RL) is well suited for. RL is a paradigm where agents learn to make decisions by performing actions in an environment and receiving feedback in the form of rewards. This approach has already led to significant achievements such as AlphaGo (Silver et al., 2016), and more recently, the success of ChatGPT (OpenAI, 2023). RL applications span various domains, including traffic systems (Chen et al., 2018), UAVs (Kaufmann et al., 2023), large language models (LLMs) (OpenAI, 2023), and dexterous manipulation (Akkaya et al., 2019).
@@
Deep RL, unlike its traditional counterparts, is typically applied in continuous state space environments, increasing the complexity of the task and thus the computational burden. In fact, with many system optimization problems, the CPU is heavily utilized by deep RL training workloads (Ali et al., 2019). However, the scalability of deep RL, particularly in learning complex state-action associations, hinges on efficiently leveraging both CPUs and GPUs (Inci et al., 2020; Espeholt et al., 2019). The combination of RL with deep neural networks necessitates a balanced approach in computational resource allocation. The processing speed, especially when updating policies involving millions of parameters, becomes a critical factor. The need to balance CPU and GPU resources, coupled with the limitations of existing frameworks, highlights the need for innovative approaches in developing efficient, scalable, and versatile systems capable of supporting the dynamic and complex nature of modern artificial intelligence and machine learning applications. This paper addresses these challenges and explores potential solutions, paving the way for more efficient and effective reinforcement learning systems.
To this end, we introduced an optimized version of Lingua Franca (LF) (Lohstroh et al., 2021), a polyglot coordination language for concurrent and time-sensitive applications. Our optimized LF is tailored to address the unique challenges of RL applications. LF stands out in its ability to effectively handle a diverse range of workloads, from lightweight, stateless computations needed for simulation to the more intensive, long-running computations required for training. A key feature of LF is its unified interface, which is adept at representing RL tasks as dataflow graphs. This offers a visual representation of the underlying RL processes, displaying the dependencies between reactors, ports, state variables, and their data. This enhances the understanding of the system’s structure, which is crucial for efficiently managing the diverse computational demands of RL applications. LF also seamlessly integrates with RL workloads, including training, simulation, and serving. On a single-node multi-core compute platform, in comparison to Ray (Moritz et al., 2018), LF achieves 1.21x and 11.62x higher simulation throughput in OpenAI Gym (Brockman et al., 2016) and Atari environments, reduces the average training time of synchronized parallel Q-learning by 31.2%, and accelerates multi-agent RL inference by 5.12x.
@@
Contributions: In this paper, we demonstrate how the reactor-oriented paradigm implemented in LF optimizes the utilization of computational resources and efficiently parallelizes RL workloads. We then describe the compilation process and the mechanisms of LF runtime, illustrating how reactors exploit parallelism and differ from the actors that underpin Ray. Our investigation also reveals that multithreading offers greater advantages than multiprocessing in parallel RL on a single node. To leverage these advantages, we introduced an optimized Reactor C runtime that supports Python without the Global Interpreter Lock, enabling true parallelism and abstracting away the burden of coordinating worker threads for RL tasks. We further present an extensive evaluation on the generations of samples from Open AI Gym and Atari Environments, synchronized parallel Q-learning, and inference of multi-agent RL. Our results demonstrate that LF outperforms Ray in terms of training, inference, and simulation for RL.
2. Parallel Reinforcement Learning and Requirements
@@
Reinforcement Learning algorithms aim to enhance an agent’s policy performance within a given environment, often represented through a simulator. These algorithms alternate cycles of data collection with the latest policy, value estimation using the latest data (and possibly data collected from previous policies), and policy improvement. Depending on the signal provided by the reward function on the collected data and the efficiency of the algorithm itself, the reinforcement learning process can vary in the amount of data and computational resources needed to run until satisfactory performance is achieved. For most tasks of interest, these data and compute requirements are large, motivating parallel reinforcement learning frameworks that can efficiently leverage multiple computational resources.
@@
Many aspects of the reinforcement learning problem can be parallelized (Samsami and Alimadad, 2020). Data collection can be split across multiple worker threads, learning can be split across multiple worker threads with each thread maintaining its own value function and policy parameters, updates to the neural networks can be performed in parallelized fashions using popular deep learning frameworks (Abadi et al., 2016; Bradbury et al., 2018; Paszke et al., 2017), and simulators can themselves be parallelized, for example on accelerated computing infrastructure (Puig et al., 2023; Szot et al., 2021; Savva et al., 2019; Shacklett et al., 2023).
2.1. Variants in RL Algorithms
@@
Single-Agent Training: The most fundamental scenario in RL is training a single agent, which involves repeatedly applying the steps of rollout, replay, and optimization. Synchronous algorithms like A2C (Mnih et al., 2016) and PPO (Schulman et al., 2017) follow these steps in sequence, whereas asynchronous variants (e.g., A3C (Mnih et al., 2016), Ape-X (Horgan et al., 2018), APPO (Zeng and Wang, 2020), IMPALA (Espeholt et al., 2018)) overlap rollout and optimization steps to enhance data throughput.
@@
Multi-Agent Training: Multi-agent training involves multiple agents interacting within the environment, either cooperatively or competitively. While the dataflow structure resembles that of single-agent training, complexities arise when customizing training for individual agents. For instance, if agents require optimization at different frequencies or are trained with distinct algorithms, the training dataflow must accommodate multiple iterative loops with varying parameters. Additionally, training populations of agents, such as Double Q-learning (Van Hasselt et al., 2015), has demonstrated great promise in RL for stabilizing training, improving exploration and asymptotic performance, and generating a diverse set of solutions.
@@
Model-Based Algorithms: Model-based RL algorithms aim to learn the transition dynamics of the environment to increase training efficiency (Moerland et al., 2023). This adds a supervised training component to parallel RL, involving training one or more dynamics models with environment-generated data.
2.2. Challenges and Opportunities in Framework Optimization:
Deep RL algorithms like Q-learning and SARSA (State-Action-Reward-State-Action) traditionally rely on sequential learning of a value table. The advent of deep neural networks has enabled these algorithms to approximate complex functions without an explicit value table. The combination of RL with deep neural networks in deep RL necessitates a balanced approach in computational resource allocation. The processing speed, especially when updating policies involving millions of parameters, becomes a critical factor. Current popular algorithms, such as Deep Q Networks (DQN) and Advantage Actor Critic (A2C), often employ a hybrid approach using both CPUs and GPUs. This setup allows for the efficient execution of different phases of the RL process, with policy evaluation typically occurring on the CPU and policy updates on the GPU. The evolution of machine learning, especially in the realm of RL, has ushered in a new era of computational requirements and challenges.
@@ 2.2. Challenges and Opportunities in Framework Optimization:
Frameworks like Ray have made strides in CPU/GPU orchestration for RL, but they face efficiency challenges due to the actor model’s inherent communication overhead and synchronization demands. Ideal frameworks for RL should support heterogeneous computations, flexibility in computational models, dynamic execution, and large data handling, while integrating seamlessly with deep learning libraries and simulation frameworks.
@@
We argue that parallel RL frameworks should better handle millions of tasks per second by providing deterministic concurrency. Current frameworks, on the other hand, fall short in efficiently meeting the evolving demands of AI applications, indicating a significant gap in the field. This research aims to propose a novel system architecture that addresses these comprehensive requirements, bridging the gap in the current landscape and pushing the boundaries of what is possible in parallel RL.
3. Introduction to Actor and Reactor Model
3.1. Actor Model
@@
The actor model (Hewitt, 2010), a foundational concept in concurrent computing, emerged in the 1970s as a response to the increasing complexity and interactivity in computer systems. It introduced a novel way to handle concurrent operations by conceptualizing ”actors” as the primary units of computation. These actors are analogous to objects in object-oriented programming (OOP) (Rentsch, 1982), but they are designed specifically to address the challenges of concurrency. Each actor represents a self-contained unit with its own local state, and actors interact with each other exclusively through asynchronous message passing. This model contrasts with traditional approaches that often rely on shared state and synchronization mechanisms like locks, which can lead to issues like deadlocks and race conditions.
@@
Furthermore, the actor model introduces a flexible approach to message processing, without enforcing any strict order in which messages must be processed. This characteristic is particularly beneficial in systems where message delivery times can vary unpredictably. Actors can process incoming messages in different sequences, and the model does not guarantee that messages will arrive in the order they were sent. This flexibility allows for more efficient utilization of resources and can lead to more robust system designs that are tolerant of delays and variable message delivery times. Overall, the actor model’s emphasis on independent, concurrent actors and asynchronous communication makes it a powerful paradigm for building scalable and resilient distributed or concurrent systems.
3.2. Reactor Model
@@
The reactor model (Lohstroh, 2020) shows great potential as an alternative to the actor model, enabling efficient and deterministic concurrency. The newly proposed reactor model represents an advancement in deterministic reactive systems, providing a structured framework for creating complex, reactive RL applications. Central to this model are the concepts of ”reactors” and ”reactions.” Reactors can be interpreted as deterministic actors, but instead of responding to messages, they react to discrete events, each linked to a specific logical time, denoted by a ”tag.” These events can trigger reactions within a reactor, similar to message handlers in actor systems, but with a key difference: reactions in reactors are governed by a defined order, ensuring determinism. Reactions are activated by discrete events, which can also be generated by reactions. Each event associates a value with a tag, representing its logical release time within the system. Reactions can access and modify state shared with other reactions in the same reactor, but interaction between different reactors is solely through events. This design choice ensures the model’s deterministic nature because the order of reaction executions is predictable and subject to strict constraints, such as tag order and execution order for reactions within the same reactor.
@@
The reactor model can be thought of as a ”sparse synchronous model” (Edwards and Hui, 2020). This means that synchronous-reactive interactions at a particular logical time may be confined to isolated parts of the system. When a reaction executes, it exclusively accesses the reactor’s state. Moreover, for reactions in the same reactor that are triggered at the same tag, their execution order is predefined, enforcing deterministic behavior. Reactors are composed of ports (inputs and outputs), hierarchy, local state, and actions. The term ”reactors” not only relate to actors in actor systems but also aligns with the synchronous reactive programming paradigm, prominent in languages like Esterel, Signal, and Lustre (Benveniste and Berry, 1991). Unlike traditional actors, reactors don’t directly reference their peers. Instead, they use named and typed ports for interconnection, enabling a hierarchical design. This hierarchy, besides facilitating deterministic behavior, also serves as a scoping mechanism for ports and imposes constraints on connection types.
@@
Additionally, reactors feature actions, a variant of ports used for scheduling future events within the same reactor or as a synchronization mechanism between internal logic and asynchronous external events. This design choice provides a bridge between deterministic internal logic and the nondeterministic external world, like sensor data, environment states, or network messages. The reactor model also incorporates state variables. The shared resources, like replay buffers, model parameters, environment states in RL, are key motivators for grouping reactions in a single reactor. However, reactors themselves do not share state, ensuring isolation between reactors and allowing parallel execution of reactions in different reactors, unless a connection necessitates sequential execution.
@@
Connections between reactors establish explicit communication channels. These connections reveal dependencies that are crucial for scheduling decisions honoring data dependencies. The reactor model simplifies the declaration of these dependencies by breaking down functionality into reactions with well-defined lexical scopes, thus eliminating dependencies out of scope. The reactor model’s execution is governed by a run-time environment. This environment is responsible for maintaining a global event queue and a reaction queue, managing logical time, and executing reactions. Details regarding the scheduler are described below.
3.3. Ray